OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 3, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (anonymous):

@zBrandz23 Please Help

OpenStudy (anonymous):

@Catch.me

OpenStudy (anonymous):

use the sequence : F(5) =F(1) + F(2) + F(4). and f(4) = f(1) + f(2) + f (3). also F(3) = f(1) + f(2) + f(2). thus compute them to get f(5).

OpenStudy (anonymous):

how do u add f1 +f2

OpenStudy (anonymous):

you have their values f(1) = 2 and f(2) = 3,

OpenStudy (anonymous):

so f4 = 5?

OpenStudy (anonymous):

@Catch.me

OpenStudy (anonymous):

how come ?? f(3) = f(1) + f(2) + F(2) = 2 + 3 + 3 = 8 then use this to compute f(4) then use f(4) to compute f(5).